OpenStudy (anonymous):

Which of the following helps you to determine that two hundred seventy-nine is divisible by 9? seventy-nine is divisible by 9. 2 plus 7 plus 9 is divisible by 9. The last digit is a 9. 9 is divisible by 9.

OpenStudy (zzr0ck3r):

if the sum of each digit is divisible by 9, then the number is divisible by 9.

OpenStudy (anonymous):

the answer is the 2nd one because 2 +7+9=18 which is divisible by 9

OpenStudy (anonymous):

so its the last one @zzr0ck3r

OpenStudy (zzr0ck3r):

well does 9 divisible by 9 implie that 2 divides 3?

OpenStudy (zzr0ck3r):

9 is of course divisible by 9, this has no affect of if 279 is divisible by 9

OpenStudy (anonymous):

no

OpenStudy (zzr0ck3r):

I said if you take the numbers 2+7+9 and get a number that is divisible by 9, then the number you started off with is divisible by 9 so which one do you think it is?

OpenStudy (anonymous):

no its the 2nd one. you decide that by seeing that all the numbers added up equalva number divisible by 3, which as you hopefully know 9 is divisible by three..... in basic terms just add up all the numbers in order to see if you can divide or not.

OpenStudy (goformit100):

@lezlezlez , @zzr0ck3r is correct at his reply he posted.

OpenStudy (zzr0ck3r):

@seasehannie we don't want them to be divisible by 3, we want them to be divisible by 9 example 12 2+1 = 3 12 is divisible by 3, but not 9.

OpenStudy (zzr0ck3r):

18 1+8 = 9 9 is divisible by 9 so 18 is divisible by 9

OpenStudy (anonymous):

what i was trying to say was that if you add up all the numbers in the number (exanmple: 27... 2+7=9 which just so happens to be divisible by 3) and see it is divisible by 3 then we can assume that the number is divisible by 9. i fully understood that we were trying to divide by 9 i just exlained it in a different way:)